# Measurability in the Fundamental Theorem of Statistical Learning

### v2

@@ A.3 Cells
$C(X)$ :=$:=\{f\colon X\to R\mid f\text{ is definable and continuous}\},$
@@
where we regard $\infty$- ∞ and $\infty$ as constant functions on $X$. For $f,g\in C_{\infty}(X)$ we write $f<g$ to indicate that $f(\underline{x})<g(\underline{x})$ for all $\underline{x}\in X$, and in this case we set
@@
    A $(0)$–cell is a singleton $\{r\}\subseteq R$ (a “point”), a $(1)$–cell is an interval $(a,b)\subseteq R$ with $a,b\in R$.
@@
We prove the statement by induction on $m$. Any $(0)$–cell is a singleton $\{r\}\subseteq R$. As $(R,\tau)$ is a Hausdorff topological space, we obtain $\{r\}\in\mathcal{B}(R)$. Any $(1)$–cell is an interval $(a,b)\subseteq R$ with $a,b\in R$, and by definition of the order-topology we obtain $(a,b)\in\tau\subseteq\mathcal{B}(R)$. In the following let $X$ be an $(i_{1},\dots,i_{m})$–cell, and assume that $X\in\mathcal{B}(R^{m})$. Further, consider $f\in C(X)$. Recall that, as a continuous function, $f$ is $\mathcal{B}(X)$$\mathcal{B}(R)$–measurable [^15] Given two measurable spaces $(\Omega_{1},\Sigma_{1}),(\Omega_{2},\Sigma_{2})$, a function $f\colon\Omega_{1}\to\Omega_{2}$ is called $\Sigma_{1}$$\Sigma_{2}$–measurable if $f^{-1}(A)\in\Sigma_{1}$ for any $A\in\Sigma_{2}$. (see Bogachev [7, Lemma 6.2.2]). In particular, the functions
@@
are $(\mathcal{B}(X)\otimes\mathcal{B}(R)$)–$\mathcal{B}(R)$–measurable. Indeed, for any set $B\in\mathcal{B}(R)$ we have $g_{f}^{-1}(B)=f^{-1}(B)\times R\in\mathcal{B}(X)\otimes\mathcal{B}(R)$ and $\pi_{2}^{-1}(B)=X\times B\in\mathcal{B}(\mathcal{X})\otimes\mathcal{B}(R)$. Thus, also the function
@@
$(g,h)_{X}$ =$=\{(\underline{x},r)\in X\times R\mid g(\underline{x})<r<h(\underline{x})\}$
=$=\{(\underline{x},r)\in X\times R\mid g(\underline{x})<r\}\cap\{(\underline{x},r)\in X\times R\mid r<h(\underline{x})\}$
@@ A.4 Examples
Consider $\mathcal{X}=\mathbb{R}$ and let $A\subseteq\mathbb{R}$ be such that $A\notin\mathcal{B}(\mathbb{R})$. Set $\mathcal{Z}=\mathbb{R}\times\{0,1\}$, $\Sigma_{\mathcal{Z}}=\mathcal{B}(\mathbb{R})\otimes\mathcal{P}(\{0,1\})$ and consider $C=A\times\{0,1\}$. Inductively applying Lemma A.10 yields $C^{m}\notin\Sigma_{\mathcal{Z}}^{m}$ for any $m\in\mathbb{N}$. Further, consider the hypothesis space =$\mathcal{H}=\{\mathds{1}_{\emptyset},\mathds{1}_{\{0\}}\}$ and the learning function $\mathcal{A}$ for $\mathcal{H}$ given by
@@
=$\mathcal{H}=\{\mathds{1}_{\emptyset}\}\cup\left\{\left.\mathds{1}_{\mathbb{R}\setminus\{w\}}\ \right|\,w\in\mathbb{R}\setminus A\right\}\!,$
@@
$\big{|}\operatorname{er}_{\mathbb{D}}(\mathds{1}_{\emptyset})-\hat{\operatorname{er}}_{\underline{z}}(\mathds{1}_{\emptyset})\hskip 0.3pt\big{|}=|1-1|=0,$
@@
$\big{|}\operatorname{er}_{\mathbb{D}}(\mathds{1}_{\mathbb{R}\setminus\{w\}})-\hat{\operatorname{er}}_{\underline{z}}(\mathds{1}_{\mathbb{R}\setminus\{w\}})\hskip 0.3pt\big{|}=|0-0|=0,$
@@
Consider $\mathcal{X}=\mathbb{R}$ and set $\mathcal{Z}=\mathbb{R}\times\{0,1\}$. As $\mathcal{B}(\mathbb{R}^{2})=\mathcal{B}(\mathbb{R})\otimes\mathcal{B}(\mathbb{R})$ and ($\mathcal{B}(\{0,1\})=\mathcal{P}(\{0,1\})$, applying Lemma A.3 yields $\mathcal{B}(\mathcal{Z})=\mathcal{B}(\mathbb{R})\otimes\mathcal{P}(\{0,1\})$. Let $\mathcal{U}(0,1)$ denote the standard uniform distribution defined on $\mathcal{B}(\mathbb{R})$, let $\delta_{1}$ denote the Dirac measure on $\{0,1\}$, and consider the distribution $\mathbb{D}=\mathcal{U}(0,1)\otimes\delta_{1}$ on $\mathcal{B}(\mathcal{Z})$. Then for any $m\in\mathbb{N}$ the probability space $(\mathcal{Z}^{m},\mathcal{B}(\mathcal{Z})^{m},\mathbb{D}^{m})$ is not complete. Indeed, for any set $A\subseteq\mathbb{R}$ with $A\notin\mathcal{B}(\mathbb{R})$, inductively applying Lemma A.10 yields $(A\times\{0\})^{m}\notin\mathcal{B}(\mathcal{Z})^{m}$. Additionally, we have $(A\times\{0\})^{m}\subseteq(\mathbb{R}\times\{0\})^{m}$, and we compute
@@
=$\mathcal{H}=\left\{\left.\mathds{1}_{\mathbb{R}\setminus\{w\}}\ \right|\,w\in\mathbb{R}\right\}\!$
@@
$\left\{\underline{z}\in\mathcal{Z}^{m}\,\left|\ U(\underline{z})\geq\tfrac{k}{m}\right.\right\}$
$=$ $\left\{\underline{z}\in\mathcal{Z}^{m}\,\left|\ \exists w\in\mathbb{R}\colon\hat{\operatorname{er}}_{\underline{z}}\big{(}h_{w}\big{)}\geq\tfrac{k}{m}\right.\right\}$
@@
$C_{w,I}^{(i)}$ =$=\begin{cases}\mathcal{Z}\setminus\Gamma\big{(}h_{w}\big{)}&\text{if }i\in I,\\ \Gamma\big{(}h_{w}\big{)}&\text{otherwise},\end{cases}$
@@
$\mathcal{R}\big{(}I,\underline{y}\>\!\big{)}\!=\!\left.\begin{cases}x_{i}=x_{j}&\text{ for any }i,j\in I\text{ with }\{y_{i},y_{j}\}=\{1\}\\ x_{i}=x_{j}&\text{ for any }i,j\in[m]\setminus I\text{ with }\{y_{i},y_{j}\}=\{0\}\\ x_{i}=x_{j}&\text{ for any }i\in I,j\in[m]\setminus I\text{ with }y_{i}=1\text{ and }y_{j}=0\\ x_{i}\neq x_{j}&\text{ for any }i,j\in I\text{ with }\{y_{i},y_{j}\}=\{0,1\}\\ x_{i}\neq x_{j}&\text{ for any }i,j\in[m]\setminus I\text{ with }\{y_{i},y_{j}\}=\{0,1\}\\ x_{i}\neq x_{j}&\text{ for any }i\in I,j\in[m]\setminus I\text{ with }\{y_{i},y_{j}\}=\{0\}\\ x_{i}\neq x_{j}&\text{ for any }i\in I,j\in[m]\setminus I\text{ with }\{y_{i},y_{j}\}=\{1\}\end{cases}\right\}.$
@@
$\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ U(\underline{z})\geq\tfrac{k}{m}\right.\right\}$

### v3

@@ A.3 Cells
$C(X)$ $:=\{f\colon X\to R\mid f\text{ is definable and continuous}\},$
@@
where we regard -$\infty$ and $\infty$ as constant functions on $X$. For $f,g\in C_{\infty}(X)$ we write $f<g$ to indicate that $f(\underline{x})<g(\underline{x})$ for all $\underline{x}\in X$, and in this case we set
@@
    A (0)(0)–cell is a singleton $\{r\}\subseteq R$ (a “point”), a (1)(1)–cell is an interval $(a,b)\subseteq R$ with $a,b\in R$.
@@
We prove the statement by induction on $m$. Any (0)(0)–cell is a singleton $\{r\}\subseteq R$. As $(R,\tau)$ is a Hausdorff topological space, we obtain $\{r\}\in\mathcal{B}(R)$. Any (1)(1)–cell is an interval $(a,b)\subseteq R$ with $a,b\in R$, and by definition of the order-topology we obtain $(a,b)\in\tau\subseteq\mathcal{B}(R)$. In the following let $X$ be an $(i_{1},\dots,i_{m})$–cell, and assume that $X\in\mathcal{B}(R^{m})$. Further, consider $f\in C(X)$. Recall that, as a continuous function, $f$ is $\mathcal{B}(X)$$\mathcal{B}(R)$–measurable [^17] Given two measurable spaces $(\Omega_{1},\Sigma_{1}),(\Omega_{2},\Sigma_{2})$, a function $f\colon\Omega_{1}\to\Omega_{2}$ is called $\Sigma_{1}$$\Sigma_{2}$–measurable if $f^{-1}(A)\in\Sigma_{1}$ for any $A\in\Sigma_{2}$. (see Bogachev [7, Lemma 6.2.2]). In particular, the functions
@@
are $(\mathcal{B}(X)\otimes\mathcal{B}(R)$$)\mathcal{B}(R$)–measurable. Indeed, for any set $B\in\mathcal{B}(R)$ we have $g_{f}^{-1}(B)=f^{-1}(B)\times R\in\mathcal{B}(X)\otimes\mathcal{B}(R)$ and $\pi_{2}^{-1}(B)=X\times B\in\mathcal{B}(\mathcal{X})\otimes\mathcal{B}(R)$. Thus, also the function
@@
$(g,h)_{X}$ $=\{(\underline{x},r)\in X\times R\mid g(\underline{x})<r<h(\underline{x})\}$
$=\{(\underline{x},r)\in X\times R\mid g(\underline{x})<r\}\cap\{(\underline{x},r)\in X\times R\mid r<h(\underline{x})\}$
@@ A.4 Examples
Consider $\mathcal{X}=\mathbb{R}$ and let $A\subseteq\mathbb{R}$ be such that $A\notin\mathcal{B}(\mathbb{R})$. Set $\mathcal{Z}=\mathbb{R}\times\{0,1\}$, $\Sigma_{\mathcal{Z}}=\mathcal{B}(\mathbb{R})\otimes\mathcal{P}(\{0,1\})$ and consider $C=A\times\{0,1\}$. Inductively applying Lemma A.10 yields $C^{m}\notin\Sigma_{\mathcal{Z}}^{m}$ for any $m\in\mathbb{N}$. Further, consider the hypothesis space $\mathcal{H}=\{\mathds{1}_{\emptyset},\mathds{1}_{\{0\}}\}$ and the learning function $\mathcal{A}$ for $\mathcal{H}$ given by
@@
$\mathcal{H}=\{\mathds{1}_{\emptyset}\}\cup\left\{\left.\mathds{1}_{\mathbb{R}\setminus\{w\}}\ \right|\,w\in\mathbb{R}\setminus A\right\}\!,$
@@
$\big|\hskip-0.3pt\operatorname{er}_{\mathbb{D}}(\mathds{1}_{\emptyset})-\hat{\operatorname{er}}_{\underline{z}}(\mathds{1}_{\emptyset})\hskip 0.3pt\big|=|1-1|=0,$
@@
$\big|\hskip-0.3pt\operatorname{er}_{\mathbb{D}}(\mathds{1}_{\mathbb{R}\setminus\{w\}})-\hat{\operatorname{er}}_{\underline{z}}(\mathds{1}_{\mathbb{R}\setminus\{w\}})\hskip 0.3pt\big|=|0-0|=0,$
@@
Consider $\mathcal{X}=\mathbb{R}$ and set $\mathcal{Z}=\mathbb{R}\times\{0,1\}$. As $\mathcal{B}(\mathbb{R}^{2})=\mathcal{B}(\mathbb{R})\otimes\mathcal{B}(\mathbb{R})$ and $\mathcal{B}(\{0,1\})=\mathcal{P}(\{0,1\})$, applying Lemma A.3 yields $\mathcal{B}(\mathcal{Z})=\mathcal{B}(\mathbb{R})\otimes\mathcal{P}(\{0,1\})$. Let $\mathcal{U}(0,1)$ denote the standard uniform distribution defined on $\mathcal{B}(\mathbb{R})$, let $\delta_{1}$ denote the Dirac measure on $\{0,1\}$, and consider the distribution $\mathbb{D}=\mathcal{U}(0,1)\otimes\delta_{1}$ on $\mathcal{B}(\mathcal{Z})$. Then for any $m\in\mathbb{N}$ the probability space $(\mathcal{Z}^{m},\mathcal{B}(\mathcal{Z})^{m},\mathbb{D}^{m})$ is not complete. Indeed, for any set $A\subseteq\mathbb{R}$ with $A\notin\mathcal{B}(\mathbb{R})$, inductively applying Lemma A.10 yields $(A\times\{0\})^{m}\notin\mathcal{B}(\mathcal{Z})^{m}$. Additionally, we have $(A\times\{0\})^{m}\subseteq(\mathbb{R}\times\{0\})^{m}$, and we compute
@@
$\mathcal{H}=\left\{\left.\mathds{1}_{\mathbb{R}\setminus\{w\}}\ \right|\,w\in\mathbb{R}\right\}\!$
@@
[…]
@@
$C_{w,I}^{(i)}$ $=\begin{cases}\mathcal{Z}\setminus\Gamma\big(h_{w}\big)&\text{if }i\in I,\\ \Gamma\big(h_{w}\big)&\text{otherwise},\end{cases}$
@@
[…]
@@
[…]